A ball is projected from the top of a tower 60m high with a velocity of 30m/s. How long does it take to reach the ground?
1
Expert's answer
2010-09-03T07:01:55-0400
We can find the time, needed for ball to reach the ground next way: V=S/t, we have S=60m and V=30m/s, t=S/V t=60/30=2s
Answer: 2s